Sawaya Co., Ltd., of Japan is a manufacturing company whose total factory overhead costs fluctuate considerably from year to year according to increases and decreases in the number of direct labor-hours worked in the factory. Total factory overhead costs at high and low levels of activity for recent years are given below:

Level of Activity

Low High
  Direct labor-hours 50,000 75,000
  Total factory overhead costs $ 14,250,000   $ 17,625,000  

The factory overhead costs above consist of indirect materials, rent, and maintenance. The company has analyzed these costs at the 50,000-hour level of activity as follows:

   
  Indirect materials (variable) $ 5,000,000
  Rent (fixed) 6,000,000
  Maintenance (mixed) 3,250,000
  Total factory overhead costs $ 14,250,000  

To have data available for planning, the company wants to break down the maintenance cost into its variable and fixed cost elements.

1.

Estimate how much of the $17,625,000 factory overhead cost at the high level of activity consists of maintenance cost. (Hint: To do this, it may be helpful to first determine how much of the $17,625,000 consists of indirect materials and rent. Think about the behavior of variable and fixed costs!)

     

2.

Using the high-low method, estimate a cost formula for maintenance.

        
       

3.

What total factory overhead costs would you expect the company to incur at an operating level of 70,000 direct labor-hours?
Answer #1

Answer to Part 1.

Indirect Materials cost at activity level of 50,000 Direct Labor Hour = $5,000,000

Indirect Materials cost per DLH = $100

Indirect Materials Cost at an activity level of 75,000 DLH = $100 * 75,000 = 7,500,000

Total Factory Overhead cost at 75,000 DLH = $17,625,000

Rent at 75,000 DLH = $6,000,000

Total Factory Overhead Cost = Indirect Materials + Rent + Maintenance

$17,625,000 = $7,500,000 + $6,000,000 + Maintenance

Maintenance at 75,000 DLH = $4,125,000

Answer to Part 2.

As per High Low method = y2 - y1/ x2 - x1

Total Maintenance cost at Highest level of Activity (y2) = $4,125,000

Total Maintenance Cost at Lowest level of Activity (y1) = $3,250,000

units at Highest level of Activity (x2) = 75,000 DLH

Units at Lowest level of Activity (x1) = 50,000 DLH

Variable cost per DLH = ($4,125,000 - $3,250,000)/ (75,000 - 50,000)

Variable Cost per DLH = $35

y = a + bx

$4,125,000 = a + ($35* 75,000)

a = $1,500,000

Cost formula for Maintenance = $1,500,000 + $35x

Answer to Part 3.

Total Maintenance Cost At 70,000 DLH = $1,500,000 + ($35 * 70,000)

Total Maintenance Cost at 70,000 DLH = $3,950,000

Total Indirect Materials cost at 70,000 DLH = $100 * 70,000 = $7,000,000

Total Factory Overhead cost at 70,000 DLH = $7,000,000 + $6,000,000 + $3,950,000

Total Factory Overhead cost at 70,000 DLH = $16,950,000